one number is 30 less that twice another number. find the 2 numbers if the sum of their squares is a minimum.
Answer by RAY100(1637)

let n=2m -30
and n^2 + m^2 = min
,
(2m-30)^2 +m^2 =0
,
4m^2 -120m +90 +m^2 =0
,
3m^2 -120 m +90 =0
3(m^2 -40m +30 ) =0
,
use quadratic eqn , a=3, b=-40,c=30
,
m= ( -(-40) +/- sqrt( -(-40) - 4 (1)(30) ) / 2(1)
m = (40 +/- sqrt (1600-120) )/2
m= (40 +/- 38.47) /2 = 39.2 ,,,, or .765
,
subst for n,
n= 2m-30 = 2*39.2 -30 = 48.4
or n= 2* .765 -30 = 28.47
,
pairs are m= 39.2, n=48.4
,
,,,,,,,,m=.765, n= 28.47
,
,
problem asks that sum of squares is a minimum, therefore use, m=.765, n=28.47